OpenStudy (anonymous):

If f(x) = 5x2 - 3x + 4 and g(x) = 2x - 5, what is f(x) • g(x) ?

OpenStudy (anonymous):

f(x)*g(x) = (5x2 - 3x + 4) ( 2x - 5) or ( 2x - 5)(5x2 - 3x + 4) = 10x^3 - 6x^2 + 8x - 25x^2 + 15x - 20 =\[\large 10x^3 - 31x^2 + 23x -20\]